# C2SaferRust: Transforming C Projects into Safer Rust with NeuroSymbolic Techniques

### v1

@@ 4.3 Implementation
We implement C2SaferRust on the nightly-2022-08-08 version of Rust. It is implemented as a cargo subcommand, so it can be run on full Cargo packages. We use the Rust compiler’s High-level Intermediate Representation (HIR) to construct the program call graph and to decompose the program into chunks. The AST traversal in Algorithm 1 is implemented as an instance of the rustc_hir::intravisit::Visitor trait. The slicing described in § 3.3 is also performed on the HIR.
We use the Mid-level Intermediate Representation (MIR) to solve the dataflow analysis for each chunk as described in § 3.3. We implement our analysis as an instance of the GenKillAnalysis trait from rustc_mir_dataflow, which computes live-in and live-out variables for a basic block. We use the MIR debug information to find the first and last basic blocks in a chunk, and get the live-in variables of the first block and the live-out variables of the last block. These become the live-in and live-out variables of the chunk.
For all our experiments, the LLM used is gpt4o-mini by OpenAI. We fixed the maximum chunk size, $L$, at 150 lines, and the number of LLM feedback iterations, $N$, at 5. In this setting, the smallest baseline took under 5 mins to translate, while the largest took nearly 14 hours.
@@ 5.1 Effectiveness of C2SaferRust in reducing unsafe constructs (RQ1)
Evaluation: We run C2SaferRust on each program in our benchmark dataset, which was compiled from coreutils. For each program, we calculated the following metrics both before and after translation (see § 4.4 for detailed information): (1) the count of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the extent of unsafe lines of code. Additionally, our analysis was expanded to encompass measurements of (1) unsafe type casts, (2) unsafe call expressions, and (3) unsafe let declarations.
@@
[…]
Before
After
$\Delta\%$
Before
After
[…]
(a) Number of raw pointer declarations and dereferences across all applications before and after translation.
[…]
@@
[…]
(b) Number of unsafe expressions and declarations AST nodes across CoreUtils projects before and after translation.
Results: § 5.1 tabulates the effectiveness of translating CoreUtils projects into safer Rust using C2SaferRust. In the following, we discuss the key findings in three key metrics: raw pointer declarations, raw pointer dereferences, and line of unsafe code.
@@
    Raw Pointer Declarations. We observe an average reduction of raw pointer declarations by $\mathbf{25\%}$ with notable reductions across all projects. The most significant improvement is observed in cat, which shows a $\mathbf{38\%}$ reduction in raw pointer declarations.
@@
We also observed a notable reduction in unsafe code (§ 5.1) across all CoreUtils projects after translation to safer Rust.
@@
    Unsafe Type Cast: Unsafe type casts decrease, but by a relatively smaller amount compared to the other metrics - an average of $\mathbf{8\%}$ reduction. The most substantial decrease is in head, from 1378 to 1189 casts (by $\approx\mathbf{14}\%$). The sub-par performance on this metric can be attributed to the extra casts needed to convert raw pointers to other types and back, e.g., 2 and 3 in Fig. 5. This is discussed in more detail below.
@@
    Unsafe Call Expression: Unsafe function calls in projects decrease by an average of 24%. The most notable reduction in calls is in uniq, from 6066 to 4397 (by $\approx\mathbf{28}\%$), highlighting the tool’s capacity to refactor and mitigate risks from low-level APIs.
@@
Discussion: Figure 5 illustrates a function using the original C code (Fig. 5), C2Rust translation (Fig. 5), and C2SaferRust’s safer output (Fig. 5) for comparison. We make the following observations:
@@
    C2SaferRust uses idiomatic and native Rust types. Non-native types in the C2Rust code (such as libc::c_char, see 1 in Fig. 5) are replaced with their native Rust types (i.e., &mut [u8] as seen in 1, Fig. 5). Note that in some cases these pointers are cast back into their raw pointer-type when passed as an argument to other functions, e.g., 2 and 3 in Fig. 5.
@@
    C2SaferRust continues to make use of FFI calls. C2SaferRust accomplishes certain core functionalities through calls to C API functions via foreign function interfaces (FFIs), e.g., safe_read and error (see 2 and 3 in Fig. 5). This limits the extent of code safety since the arguments and return types of these functions still use raw pointer types. These may only be eliminated when all the external functions are replaced with safe Rust equivalents; this is beyond the scope of this work.
@@
    C2SaferRust eliminates nonstandard control flow structures and dead code. Rust code generated by the C2Rust transpiler generates a large number of unreachable branches in the C2Rust code. It makes explicit calls to branch conditions that cannot be executed (e.g., if 0 != 0 as seen in 4, Fig. 5). C2Rust also uses certain macros such as unreachable!() (as seen in 5, Fig. 5). C2SaferRust effectively eliminates these dead code fragments and produces more concise and idiomatic Rust code.
@@
Summary: C2SaferRust reduces unsafe constructs in CoreUtils, cutting raw pointers by 38% and dereferences by 27%. Unsafe code lines, type casts and call expressions decrease by 24%, 8% and 24%, enhancing Rust safety and idiomaticity.
@@
Figure 6: For our coreutils benchmark dataset - a plot of the fraction of successful translations versus function length, and the number of attempts required to translate a function versus function length. This includes only functions that are translated as a single unit, i.e., functions that are under 150 lines.
Results: The results for the component ablation are shown in Table IV. There are 3 utilities and 5 metrics, which makes a total of 15 settings. C2SaferRust outperforms both “No Chunking” as well as “Random Order” in 9 out of 15 settings, showing a clear benefit of both design choices.
LABEL:fig:success_vs_length shows the rate of translation success versus function length, and Figure 5(b) shows the number of attempts needed to translate a function vs its length. If we ignore the outliers to the right of the distribution, we can see that as the function length increases, the translation success rate decreases and the number of required attempts increases. This clearly indicates that the longer a function is, the harder it is for an LLM to translate it. This provides empirical support for our decision to break up very long functions into smaller chunks for translation.
@@ 6 Related Work
Converting C to unsafe Rust. Corrode [6] and C2Rust [10] are two systems that aim to convert C code to functionally equivalent Rust code. Both tools have limitations and corner cases that they cannot handle. The primary difference between them is that C2Rust is still being actively developed and maintained, while Corrode has not been updated for several years. The Rust code that these tools produce is non-idiomatic and makes extensive use of unsafe features. Citrus [5] is a third tool that uses syntactic mappings to convert C to non-idiomatic, unsafe Rust. However, unlike Corrode and C2Rust, Citrus doesn’t aim to produce functionally equivalent code. The translated code may not run or compile. Citrus aims to be a lighter alternative to C2Rust, to produce a first draft of Rust code that can manually be refactored into a working Rust program.
Converting unsafe Rust to safer Rust. References in Rust are tracked by the borrow checker to ensure that they don’t extend beyond the scope of the value they are pointing to. However, the non-idiomatic Rust code generated by C2Rust makes extensive use of raw pointers, which are not tracked by the borrow checker. There has been some work on automatically converting a subset of the raw pointers in a program to references, thus creating safer Rust code.
Laertes [8] was the first such approach, which used compiler errors to extract information about lifetimes and mutability of each raw pointer, and tried to iteratively rewrite each one as a borrow. The authors also performed an in-depth study about how the different ways in which unsafe features are used in C2Rust transpiled code. However, this method can only be applied to functions where the cause of unsafety is due to a lack of ownership information on a raw pointer. Only 20% of the functions fall into this category, and they are only able to convert about 8.3% of the raw pointers in their benchmark dataset to references. In a follow-up paper [9], the authors of Laertes try to expand their technique to include more categories of raw pointers. With this new approach, they are able to convert about 21% of the pointers to references.
CROWN [30] is another approach that starts with C2Rust code and tries to convert each raw pointer into a reference. It derives constraints for each raw pointer, and solves these constraints using an SMT solver to convert each raw pointer into a reference. However, like Laertus, CROWN too has a limitation - it can handle only mutable, non-array raw pointers. It succeeds at converting about 37% of these to references.
Hong and Ryu [11] develop a tool to perform a very specific type of transformation on C2Rust code - replacing output parameters that indicate success/failure with canonical Rust types like Option and Result. In their subsequent work [12], they study another specific type of transformation on C2Rust code - replacing C-style unions with Rust tagged unions.
LLMs for Code Translation. Roziere et al. [21] proposed TransCoder, an early attempt to tackle code translation using transformer models. They proposed Back Translation, a fully unsupervised approach. Ahmad et al. [2] refined Back Translation with the addition of a code summarization task. In contrast to unsupervised learning, supervised learning requires aligned parallel data across pairs of programming languages, which is challenging to obtain at scale. Roziere et al. [22] addressed this challenge by generating synthetic aligned data using another language model and filtering out incorrect translations using generated unit tests. The last few years have seen the development of LLMs for code and natural language, that are pretrained on massive amounts of data with an autoregressive objective function. Pan et al. [19] and Yang et al. [29] perform thorough empirical evaluations of the translation ability of these models. Yang et al. [29] also propose UniTrans, a tool that generates test cases, provides them along with the code to an LLM for translation, and uses test feedback to iteratively repair the generated translation. Nitin et al. [18] propose to generate multi-modal specifications with a self-consistency filter and provide these along with the code to an LLM for translation. [28] propose to verify candidate LLM translations against the output of a transpiler using rigorous equivalence proving.
However, all of these approaches deal only with settings where the entire program is relatively short and fits within a single LLM context window. The typical setting is that of solutions to competitive programming problems. Shiraishi and Shinagawa [25] was the first work to attempt translation of arbitrarily large programs using LLMs. They slice a C repository into smaller translation units, translate each unit to Rust using an LLM, and use compiler feedback to fix translations. However, they do not use test feedback, and they require each translation unit to be independently compilable.
There have been a few prior works that attempt to repository-level translation for other pairs of languages. Ibrahimzada et al. [13] attempt translation of full Java projects to Python. Zhang et al. [31] translate Go to Rust, focusing on avoiding runtime errors in the translated code. While preparing this manuscript, we came across a recent ArXiv preprint [24] that presents an approach to perform full repository C to Rust translation. They divide the program into translation units, translate each unit using an LLM, and use dynamic analysis along with LLM-based test generation to verify the correctness of each translation. However, our approach is different in certain important aspects. In particular, a) we use a transpiler to provide a skeleton for the target language code, thus allowing us to check end-to-end correctness at every step; b) we break down code up to a sub-method level, ensuring that we can scale to programs with methods of arbitrary length.

### v2

@@ 4.3 Implementation
We implement C2SaferRust on the nightly-2022-08-08 version of Rust. It is implemented as a cargo subcommand, so it can be run on full Cargo packages. We use the Rust compiler’s High-level Intermediate Representation (HIR) to construct the program call graph and to decompose the program into chunks. The AST traversal in Algorithm˜1 is implemented as an instance of the rustc_hir::intravisit::Visitor trait. The slicing described in §˜3.3 is also performed on the HIR.
We use the Mid-level Intermediate Representation (MIR) to solve the dataflow analysis for each chunk as described in §˜3.3. We use the MIR debug information to find the first and last basic blocks in a chunk, and get the live-in variables of the first block and the live-out variables of the last block. These become the live-in and live-out variables of the chunk.
For all our experiments, the LLM used is gpt4o-mini by OpenAI. We fixed the maximum chunk size, $L$, at 150 lines, and the number of LLM feedback iterations, $N$, at 5. The smallest baseline, qsort, took under 5 mins to translate, and the largest, optipng, took nearly 14 hours. More details about LLM calls and costs are in the Appendix.
@@ 5.1 Effectiveness of C2SaferRust in reducing unsafe constructs (RQ1)
Evaluation: We run C2SaferRust on each program in our benchmark dataset, which was compiled from coreutils. For each program, we calculated the following metrics both before and after translation (see §˜4.4 for detailed information): (1) the count of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the extent of unsafe lines of code. Additionally, our analysis was expanded to encompass measurements of (1) unsafe type casts and (2) unsafe call expressions.
@@
    Raw Pointer     Raw Pointer     Unsafe     Unsafe Call       Unsafe
    Declarations ($\nabla$)     Deferences ($\nabla$)     Lines ($\nabla$)     Expressions ($\nabla$)     Type Casts ($\nabla$)
@@
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     252    214    15    656    540    18    11324    9250    18    2353    2212    6    5979    4663    22
   pwd     164    129    21    295    225    24    4201    3151    25    875    871    0    2248    1563    30
   cat     192    120    38    317    240    24    5625    4189    26    1038    912    12    3116    1992    36
   truncate     156    124    21    326    263    19    5544    4443    20    1040    952    8    3357    2521    25
   uniq     227    166    27    343    250    27    6066    4397    28    1150    1025    11    3590    2340    35
   tail     389    297    24    1092    847    22    11663    8818    24    2580    2433    6    5869    3739    36
   head     192    141    27    442    351    21    6245    4699    25    1378    1189    14    3488    2464    29
     avg. $\Delta\%$             25            22            24            8            30
(a) Our metrics for measuring safety of Rust code, computed across all applications before and after translation.
    Raw Pointer     Raw Pointer     Unsafe     Unsafe Call     Unsafe
    Declarations ($\nabla$)     Deferences ($\nabla$)     Lines ($\nabla$)     Expressions ($\nabla$)     Type Casts ($\nabla$)
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
Before
After
$\Delta\%$
     split     141    130    8    463    405    13    6836    6134    10    1405    1334    5    5400    4427    18
   pwd     38    40    -5    85    74    13    704    565    20    217    223    -3    646    292    55
   cat     93    56    40    242    200    17    3508    $28$    19    659    572    13    2625    1587    40
   truncate     68    66    3    241    211    12    4050    $36$    10    772    676    12    4231    3332    21
   uniq     85    71    16    156    144    8    3027    2636    13    569    516    9    3377    2917    14
   tail     147    125    15    568    475    16    7203    6191    14    1559    1514    3    5619    4129    27
   head     87    75    14    340    297    13    3724    3160    15    903    801    11    3193    2355    26
     avg. $\Delta\%$             13            13            14            7            29
(b) Our metrics for measuring safety of Rust code, computed across all applications, filtered to include covered functions only.
Results: Table˜II tabulates the effectiveness of translating CoreUtils projects into safer Rust using C2SaferRust. In the following, we discuss the key findings in three key metrics: raw pointer declarations, raw pointer dereferences, and line of unsafe code.
@@
    Raw Pointer Declarations. We observe an average reduction of raw pointer declarations by $\mathbf{25\%}$ with notable reductions across all projects. The most significant improvement is observed in cat, which shows a $\mathbf{38\%}$ reduction in raw pointer declarations. Unexpectedly, we see an increase in this metric for pwd in the filtered seting. This does not signify a decrease in safety of the program; rather, it is due to a quirk of the way this metric is computed. The metric counts only variable bindings as declarations. So foo("abc" as const *u8) would not be counted as a raw pointer declaration, but let arg = "abc" as const *u8; foo(arg) would be.
@@
We also observed a notable reduction in unsafe code across all CoreUtils projects after translation to safer Rust.
@@
    Unsafe Call Expression: Unsafe function calls decrease, but by a relatively smaller amount compared to the other metrics - an average of $\mathbf{8\%}$ reduction. The most substantial decrease is in head, from 1378 to 1189 casts (by $\approx\mathbf{14}\%$). The sub-par overall performance is due to the reliance on FFI calls, which is discussed further below.
@@
    Unsafe Type Cast: Unsafe type casts decrease by an average of 30%. The most substantial decrease is in tail, from 5869 to 3739 casts (by $\approx\mathbf{36}\%$).
@@
Discussion: Figure˜5 illustrates a function using the original C code (Fig.˜5), C2Rust translation (Fig.˜5), and C2SaferRust’s safer output (Fig.˜5) for comparison. We make the following observations:
@@
    C2SaferRust uses idiomatic and native Rust types. Non-native types in the C2Rust code (such as libc::c_char, see 1 in Fig.˜5) are replaced with their native Rust types (i.e., &mut [u8] as seen in 1, Fig.˜5). Note that in some cases these pointers are cast back into their raw pointer-type when passed as an argument to other functions, e.g., 2 and 3 in Fig.˜5.
@@
    C2SaferRust continues to make use of FFI calls. C2SaferRust accomplishes certain core functionalities through calls to C API functions via foreign function interfaces (FFIs), e.g., safe_read and error (see 2 and 3 in Fig.˜5). This limits the extent of code safety since the arguments and return types of these functions still use raw pointer types. These may only be eliminated when all the external functions are replaced with safe Rust equivalents; this is beyond the scope of this work.
@@
    C2SaferRust eliminates nonstandard control flow structures and dead code. Rust code generated by the C2Rust transpiler generates a large number of unreachable branches in the C2Rust code. It makes explicit calls to branch conditions that cannot be executed (e.g., if 0 != 0 as seen in 4, Fig.˜5). C2Rust also uses certain macros such as unreachable!() (as seen in 5, Fig.˜5). C2SaferRust effectively eliminates these dead code fragments and produces more concise and idiomatic Rust code.
@@
Summary: C2SaferRust reduces unsafe constructs in Coreutils, cutting raw pointers by 38% and dereferences by 27%. Unsafe lines, type casts and calls decrease by 24%, 8% and 24%, enhancing safety and idiomaticity.
TABLE III: A comparative evaluation of C2SaferRust and other state-of-the-art methods against the baseline (C2Rust). Lower values are preferred. Cells in blue show the greatest improvement on that metric for that project. C2SaferRust presents the most significant improvement as indicated by the best Win/Tie/Loss ratio. Laertes and crown actually cause an increase in the amount of unsafe code in several translations, as indicated by the negative values under the $\%\Delta$ columns.
Raw Pointer
Declaration
Raw Pointer
Dereferences
Lines of
Unsafe Code
Unsafe Call
Expressions
Unsafe
Type Casts
    $\Delta$    $\Delta$    $\Delta$    $\Delta$    $\Delta$
      bzip2     Baseline     227     3764     12594     2580     7614
        Laertes     206 9    3699 2    13115 -4    2788 -8    8699 -14
        crown     194 15    3674 2    12592 0    $29$ -14    7672 -1
        C2SaferRust     199 12    2691 29    10328 18    2455 5    4097 46
      genann     Baseline     73     339     1634     579     1319
        Laertes     73 0    339 0    1650 -1    581 -0    1329 -1
        crown     71 3    317 6    1775 -9    643 -11    1317 0
        C2SaferRust     63 14    273 19    994 39    578 0    362 73
      lil     Baseline     438     1668     4852     1809     2788
        Laertes     430 2    1658 1    4645 4    1844 -2    $27$ 2
        crown     355 19    756 55    5073 -5    3383 -87    2523 10
        C2SaferRust     428 2    1449 13    4582 6    1843 -2    2111 24
      tulipindicators     Baseline     866     1847     17671     2595     12983
        Laertes     866 0    1847 0    25594 -45    2595 0    13050 -1
        crown     865 0    1847 0    17769 -1    2419 7    13195 -2
        C2SaferRust     860 1    1625 12    17314 2    2638 -2    12274 5
      urlparser     Baseline     79     60     845     447     648
        Laertes     74 6    2 97    1002 -19    572 -28    1167 -80
        C2SaferRust     81 -3    51 15    702 17    450 -1    439 32
      grabc     Baseline     13     21     214     48     123
        Laertes     8 38    12 43    213 0    71 -48    122 1
        C2SaferRust     7 46    11 48    84 61    33 31    10 92
      optipng     Baseline     1407     6062     82616     7854     32913
        Laertes     1235 12    5552 8    93696 -13    9189 -17    46450 -41
        C2SaferRust     1197 15    5203 14    64013 23    7354 6    25061 24
      qsort     Baseline     4     10     32     11     12
        Laertes     2 50    6 40    32 0    23 -109    12 0
        C2SaferRust     1 75    6 40    19 41    10 9    13 -8
      snudown     Baseline     244     842     5842     1587     3252
        Laertes     224 8    765 9    7213 -23    1772 -12    5940 -83
        C2SaferRust     231 5    747 11    5583 4    1533 3    2867 12
      xzoom     Baseline     29     172     1468     321     867
        Laertes     29 0    172 0    1530 -4    321 0    897 -3
        C2SaferRust     29 0    168 2    1467 0    325 -1    631 27
Aggregate
Win/Tie/Loss
C2SaferRust     5/1/4     7/1/2     10/0/0     8/0/2     9/0/1
        crown     2/0/8     1/0/9     0/0/10     1/0/9     0/0/10
        Laertes     1/1/8     1/1/8     0/0/10     1/0/9     1/0/9
5.2 Comparison with Existing Tools (RQ2)
Motivation: This section presents a comparative analysis of safe Rust transformations obtained from C2SaferRust, as well as methodologies from contemporary approaches Laertes and crown applied to Laertes benchmark. [^5] The incompatibility of Laertes with the current version of the Rust compiler precluded its execution on our CoreUtils benchmarks due to it being developed on a deprecated compiler version. Therefore, C2SaferRust was instead executed on the 10 Laertes benchmarks666crown and was executed on 4 of the 10 benchmark projects. In the remaining projects, we encountered failure due to limitations in managing unions and variadic arguments [30].
Evaluation: We measure heuristics similar to the previous research question: (1) the number of raw pointer declarations, (2) the number of raw pointer dereferences, and (3) the quantity of unsafe lines of code. Furthermore, we extend our analysis to include measurements of (1) unsafe type casts, (2) unsafe call expressions, and (3) unsafe let declarations.
Results: The results are shown in Table˜III. We observe that C2SaferRust is capable of reducing the size of unsafe blocks, doing better than the baselines in most cases.
  • On the two raw pointer metrics (i.e., raw pointer declarations and raw pointer dereferences), C2SaferRust performs as well as or better than the baselines on a majority of the benchmark programs with an aggregate Win/Tie/Loss of 5/1/4 and 7/1/2 for raw pointer declarations and raw pointer dereferences respectively.
  • C2SaferRust demonstrates superior efficacy in reducing lines of unsafe code, achieving a reduction of up to 61% in grabc. An aggregate Win/Tie/Loss of 10/0/0 shows consistent performance across all applications.
  • We also observed a notable reduction in key unsafe expressions and declarations with C2SaferRust in all projects. On the other hand, Laertes and crown actually cause an increase in unsafe code in several projects, as evidenced by the negative values of $\%\Delta$ in Table˜III under lines of unsafe code, unsafe type casts and unsafe calls.
Section E in the Appendix shows an example of the outputs of C2SaferRust and Laertes, for qualitative comparison.
{myshadowbox}
Summary: C2SaferRust outperforms other approaches with a 39/2/9 aggregate win/tie/loss record over all 5 metrics, enhancing Rust safety and idiomaticity.
5.3 Ablation of C2SaferRust design choices (RQ3)
In this section, we evaluate the individual effect of some of the components of C2SaferRust.
Evaluation: We pick 3 arbitrary baselines from coreutils - cat, uniq and head - and run C2SaferRust in two settings:
  1. 1.
    No Chunking: Translating each function as a whole without dividing it into smaller chunks. Our hypothesis is that LLMs do better when the input is not too long, so we would expect this setting to be worse than the original.
  2. 2.
    Random Order: Translating the functions in a random order, instead of following the ordering described in §˜3.2.2.
Function Length: We also study the effect of function length on translation. For this study, we consider only those functions which are translated as a single unit, i.e., those which are smaller than $L=150$ lines. We define two quantities:
  • Translation Success Rate: The percentage of functions that translate correctly, where a correct translation is one that compiles and passes tests.
  • Number of Attempts: The number of iterations of compiler or test feedback given to the LLM before it generates a successful translation. This is measured only for functions for which a correct translation is generated within $N$ attempts.
We measure Translation Success Rate and Number of Attempts over functions of varying length, from 0 to $L=150$.
(a) Translation Success Rate versus Function Length
@@ 5.3 Ablation of C2SaferRust design choices (RQ3)
Figure 6: For our coreutils benchmark dataset - a plot of the fraction of successful translations versus function length, and the number of attempts required to translate a function versus function length. This includes only functions that are translated as a single unit, i.e., functions that are $\leq$ 150 lines.
Results: The results for the component ablation are shown in Table˜IV. There are 3 utilities and 5 metrics, which makes a total of 15 settings. C2SaferRust outperforms both “No Chunking” as well as “Random Order” in 9 out of 15 settings, showing a clear benefit of both design choices.
Figure˜6(a) shows the rate of translation success versus function length, and Figure˜6(b) shows the number of attempts needed to translate a function vs its length. If we ignore the outliers on the right, we can see that as the function length increases, the translation success rate decreases and the number of attempts increases. This provides empirical support for our decision to break up very long functions into smaller chunks for translation.
@@ 6 Related Work
Converting C to unsafe Rust. Corrode [5] and C2Rust [9] both aim to convert C to functionally equivalent Rust. C2Rust continues to be actively maintained, whereas Corrode has seen little recent development. Both generate non-idiomatic Rust that relies heavily on unsafe constructs. Citrus [4] similarly provides syntactic mappings from C to Rust, but focuses on being lightweight rather than fully correct: the output may not compile or run without manual revision. It is intended more as a rough starting point than a complete translation pipeline.
Converting unsafe Rust to safer Rust. Because C2Rust code relies extensively on raw pointers, much work has explored automatically replacing them with safer Rust references. Laertes [7] was the first to use compiler feedback to infer lifetimes and mutability, rewriting raw pointers when ownership information could be recovered. Their study also analyzed common patterns of unsafe in C2Rust output. However, the approach applies only when unsafety arises from missing ownership rather than aliasing or other causes, converting about 8.3% of raw pointers. A later extension [8] broadens coverage to more pointer categories and improves conversion rates to about 21%. CROWN [30] also starts from C2Rust output but instead constructs pointer constraints and solves them using an SMT solver. It is more targeted, supporting only mutable, non-array pointers, and converts $\sim$37% of these to safe references. Other efforts pursue narrower correctness improvements: Hong and Ryu [10] replace output-parameter success/failure patterns with Option/Result, and [11] translate C unions into Rust tagged unions. NADER [19] replaces unchecked indexing with checked indexing, allowing a tunable safety–performance tradeoff, but does not address raw pointers or unsafe blocks more generally.
LLMs for Code Translation. Early LLM-based translation work began with TransCoder [21], using Back Translation for fully unsupervised learning. Later, Ahmad et al. [2] enhanced this with code summarization. Roziere et al. [22] generate aligned parallel data for supervised training by generate synthetic pairs using another model and verifying them with unit tests. Recent code-focused LLMs have demonstrated strong translation capabilities [17]. Yang et al. [29] introduce UniTrans, which generates tests and iteratively repairs translations using execution feedback. Nitin et al. [16] use multi-modal specifications with a self-consistency filter, and Yang et al. [28] verify candidate translations against transpiler output using equivalence proofs. However, these works generally deal with programs that fit in a single model context window, such as competitive programming solutions.
Shiraishi and Shinagawa [25] were the first to attempt translation of arbitrarily large repositories, slicing projects into smaller units, translating each with an LLM, then compiling and repairing errors. Their method does not incorporate test feedback and requires each unit to compile independently, limiting scalability. Ibrahimzada et al. [12] translate full Java projects to Python, and Zhang et al. [31] translate Go to Rust with emphasis on avoiding runtime errors. Shetty et al. [24] perform full C→Rust repository translation by dividing code into units, translating with LLMs, and validating using dynamic analysis with LLM-generated tests. Our approach differs in two key ways: (a) we leverage a transpiler to generate Rust skeletons, enabling end-to-end correctness checking at every iteration; and (b) we decompose code below method granularity, allowing translation of repositories with extremely large functions.